In a storehouse, 80 goods are kept where 20 of them are damaged. If a good is taken in random, what is the probability of non-damaged good is taken?
How can I solve this? I've tried using probability of c(80,60) but I still can't find the answer.
The choices are:
A. 1/8 B. 3/4 C.1/10 D. 1/20
Thank you

Answer by ramkikk66(644)


Hi,

C(80,60) is the number of ways you can choose a set of 60 from a set of 80. But
that's not what you want here. Here, you are choosing ONE item from 80, because
the problem states that you are choosing one good at random from 80. This 
can be done in C(80,1) or 80 ways. Similarly, there are 60 non-damaged goods and 
you can choose one from them in C(60,1) = 60 ways.

Probability is defined as Number of favourable outcomes / No. of total possible outcomes.

Here the probability = No. of ways you can choose a non-damaged good / No. of
ways of choosing any one good

= C(60,1)/C(80,1) = 60/80 = 3/4

Hope you got it. Please email me if it's not clear.

:)